OpenStudy (anonymous):

A can of soda at 83° F. is placed in a refrigerator that maintains a constant temperature of 37° F. The temperature T of the soda t minutes after it is placed in the refrigerator is given by T(t) = 37 + 46e – 0.058 t Find the temperature of the soda 10 minutes after it is placed in the refrigerator. (Round to the nearest tenth of a degree.)

OpenStudy (anonymous):

\[T(0) = 37 + 46e - 0.058\times0\] \[T(t) = T(0) - 0.058 \times t\] \[T(10) = T(0) - 0.58\] \[T(10) = 83 - 0.58 = 82.42 \approx 82.4\]

OpenStudy (datanewb):

This question does not involve linear algebra techniques and the equation for T(t) is wrong. It should look more like this: T(t) = Ts + (To - Ts)*e^(-k*t) Where k is given as .058, t = 10, To = 83, and Ts = 37. T(10) = 37 + (83 - 37)*e^(-.058 * 10) \[T(10) \approx62.8\] In any event, this question should be closed/ moved to the appropriate forum. Please post Linear Algebra related questions here. Thank you.